OpenStudy (genericnoodles):

I have an Algebra 1 problem in stuck on if someone could help :) If ab=37, and =1/37, what is the value of the product x•b•y•a? I tried it and I got 1, but if someone could explain it further It'd be awesome.

OpenStudy (anonymous):

you forgot to put what 1/37 equals.

OpenStudy (anonymous):

Guessing the typo is xy=1/37 .

OpenStudy (genericnoodles):

No, that's the whole problem. I'm rather confused, but shouldn't xy= 37/1?

OpenStudy (anonymous):

Are you sure it's just =1/37 in the middle of the sentence?

OpenStudy (anonymous):

lol

OpenStudy (genericnoodles):

I'm typed the question exactly like it is in the book. This is in the commutative, distribution, and associative part of the book, I'm not sure how to apply those to the problem, but yes I didn't leave anything out.

OpenStudy (genericnoodles):

I had to do this today in class and the other students couldn't figure it out either. Maybe it is a miss type, I just want to be sure first.